For this problem, carry at least four digits after the decimal in your calculations. Answers may vary slightly due to rounding.

In a random sample of 69 professional actors, it was found that 35 were extroverts.

(a) Let p represent the proportion of all actors who are extroverts. Find a point estimate for p. (Round your answer to four decimal places.)


(b) Find a 95% confidence interval for p. (Round your answers to two decimal places.)

lower limit    
upper limit    


Give a brief interpretation of the meaning of the confidence interval you have found.

We are 5% confident that the true proportion of actors who are extroverts falls within this interval.We are 95% confident that the true proportion of actors who are extroverts falls outside this interval.    We are 95% confident that the true proportion of actors who are extroverts falls within this interval.We are 5% confident that the true proportion of actors who are extroverts falls above this interval.


(c) Do you think the conditions np > 5 and nq > 5 are satisfied in this problem? Explain why this would be an important consideration.

Yes, the conditions are satisfied. This is important because it allows us to say that is approximately normal.No, the conditions are not satisfied. This is important because it allows us to say that is approximately binomial.    No, the conditions are not satisfied. This is important because it allows us to say that is approximately normal.Yes, the conditions are satisfied. This is important because it allows us to say that is approximately binomial.
Answer #1

(a) The point estimate \hat{p} = 35/69 = 0.5072

(b) The Zc for \alpha = 0.05 is 1.96. \hat{p} = 0.5072 and 1 - \hat{p} = 0.4928.

The Margin of error ME = 1.96 * Sqrt(0.5072*0.4928/69) = 0.1180

The Lower Limit = \hat{p} - ME = 0.5072 - 0.1180 = 0.3892 \approx0.39

The Upper Limit = \hat{p} + ME = 0.5072 + 0.1180 = 0.6252 \approx0.63

We are 95% confident that the true proportion of actors who are extroverts falls within this interval.

(c) Yes, the conditions are satisfied. This is important because it allows us to say that \hat{p} is approximately normal.